OpenStudy (anonymous):

6,17,28,39,50 find the formula for its general term.

OpenStudy (hoblos):

the difference between any two consecutive numbers is 11 so the formula of general term is U(n) = U(n-1) +11

OpenStudy (anonymous):

its un=11n-5

OpenStudy (hoblos):

yeah if you want it in terms of n it would be Un = 11n-5 but also it is an arithmetic sequence that can be written as Un = U(n-1) +11 where U(1) = 6

OpenStudy (anonymous):

it is an arithmetic sequence where a (z 1st term) is 6. the common difference is 11 so a+(n-1)d simply and replace to get 6+(n-1)11 6+11n-11 11n-5 (answer)